# Topological excitations in a spin-orbit coupled spin-1 Bose-Einstein condensate under sinusoidally varying magnetic fields

### v1

@@ 3.1 Polarized gradient magnetic field
Figure 1: The ground state densities and momentum distributions for the spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC. The first and second columns show the density ($a1-a3$) and momentum ($b1-b3$) distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components for the $x-$ polarised magnetic field $B=B_{0}x$. The third and fourth columns represent the density ($c1-c3$) and momentum ($d1-d3$) distributions for the $y-$ polarised magnetic field $B=B_{0}y$. The magnetic field strength $B_{0}=0.25$ for both cases. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
Now, with SOC in play, our investigation centers to unravel the competition between the SOC and external magnetic field. Depending on the choices of the external magnetic field applied, the ground state structures for the SO-coupled spin-1 BEC are different. We show the ground state structures for the SOC strength $\gamma=0.5$ for the $x-$polarised magnetic field in figure 2. The ground state densities $n_{\pm 1},n_{0}$ exhibit modulations along the $x=0$ line as shown in Fig. 2 ($a1-a3$). Our numerical simulations show that the formation of vortices in the three components is pertinent to the interplay of SOC and the external magnetic field. The SOC interacts with the magnetic field in such a way that the effective magnetic field is modified. In addition to this, as the SOC couples the intrinsic spin to the particles’ motion, it imparts some extra angular momentum to the system which results in the formation of vortices in the system. Although there are no saddle points in the magnetic field, these vortices appear at the $x=0$
@@
Figure 2: The ground state structures for the spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the the $x-$ polarised magnetic field $B=B_{0}x$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
@@
Figure 3: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 2. The unit length is the oscillator length $l$.
@@
line under the influence of SOC. The vorticity profile associated to each component is shown in Fig. 2 ($b1-b3$). The vortices are slightly deformed along the $y-$direction as the component densities are modulated along the $x=0$ line. The k-space density profiles (as shown in Fig. 2 (c1-c3)) depict that apart from the central peaks, some extra momenta predominantly along the $k_{x}-$direction is observed. These extra peaks at $k_{x}\neq 0$ corresponds to the density modulations along the $x-$ direction. This is due to the fact that SOC influences the distribution of particles in the k-space, similar to the impact of SOC as seen in the distribution of electronic states as a function of momentum in solid-state physics [71].
@@
The combined effect of the in-plane direction dependent gradient field and the SOC yield intricate spin textures in the spin-1 BEC as shown in Fig. 3 ($b$) with corresponding topological charge distribution $q$ shown in Fig. 3($a$). We can identify the presence of a skyrmion at the center of the vortex core, from the spin texture. The local amplification of the skyrmion spin texture is shown in Fig. 3 ($c$).
For the SO-coupled spin-1 BEC in $y-$polarised gradient magnetic field, we can analogously understand the density modulations along the $y=0$ line for all the three components. However, these density modulations converge into the formation of anti-vortices in the system, and the corresponding spin textures have anti-skyrmions along the $x-$axis. It is interesting to note that the state which had fragments in the absence of SOC, now has more vortices or anti-vortices as we observe for the case of $\gamma=0.5$. As we increase the SOC strength $\gamma$, the number of vortices increases. However, for the polarised magnetic field, they appear in the same alignment, with the vortex spacing being reduced.
3.2 Cross-polarised sinusoidally varying in-plane magnetic fields
Thus far, we have discussed how impactful the polarised gradient field can be on the ground state structure of a SO-coupled spin-1 BEC. To further exploit this direction-dependent interplay of external magnetic field and SOC, we now choose sinusoidally varying in-plane magnetic fields in both $x$ and $y$ directions. In general, to make the magnetic field divergenceless, we also apply a magnetic field $B_{z}=-B_{0}zf(x,y)$ in the $z$-direction, where $f(x,y)$ are sinusoidal functions in the $x-y$ plane. However, as we are in the Quasi-2D geometry, for tight binding along the $z-$direction, $B_{z}$ will not affect the spinor condensate. In this section, we center our focus only on the in-plane sinusoidal magnetic field. Experimentally, a polarised sinusoidal field can be realised using a similar technique discussed in [18], shown for the magnetic field in the vertical direction. Campbell et al has also demonstrated a similar in-plane sinusoidal field in their work [72]. We discuss below, various combinations of $B(\mathbf{r})=(B_{x},B_{y})$ in different cases.
3.2.1 Case-\@slowromancapi@ Sine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{sin}y)$
For the present case, we choose the magnetic fields as sine functions in both directions, i.e., $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with equal magnetic strength $B_{0}=0.25$ in both directions. We first present the ground state structures for the no SOC case in Fig. 4. The density profiles $n_{1},n_{0}$ and $n_{-1}$ of the condensates $m_{F}=1,m_{F}=0$ and $m_{F}=-1$, respectively are shown in Fig. 4 ($a1-a3$). The sine-sine type of magnetic field results in the density modulations in $m_{F}=\pm 1$ components with the appearance of topological exciations in the form of a triad of alternating vortex-anti-vortex along both $x-$ and $y-$axes. This is due to the fact that the applied magnetic field from both directions, generate certain saddlepoint structures such that the nature of the triads in $m_{F}=1$ and $m_{F}=-1$ components are complementary to each other. The opposite nature of the triads is due to the fact that the magnetic field appears as conjugate for the $m_{F}=1$ and $m_{F}=-1$ components in the GP equations (2-2). Furthermore, combining with the spin-mixing term, the magnetic field yields vortex brights in the $m_{F}=0$ component at
@@ 3.2.1 Case-\@slowromancapi@ Sine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{sin}y)$
Figure 4: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the sine-sine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
the loci (saddle points of $B(\mathbf{r})$) of vortices in the $m_{F}=\pm 1$ components so as to conserve the overall angular momentum of the spin-1 BEC in the absence of SOC. Thus, the $m_{F}=0$ component has no vortices, instead a high density peak is observed at the center of the condensate. The vorticity profiles associated to the condensate densities are depicted in Fig. 4 ($b1-b3$). These vortices or anti-vortices are coreless as there is no phase singularity in the total density of the system. The k-space density profiles (see Fig. 4 ($c1-c3$)) support observed density modulations in the system. For the $m_{F}=1$ and $m_{F}=-1$ components, apart from the central peak, we observe finite elongations in both the $k_{x}$ and $k_{y}$ directions which correspond to the density modulations with wavelength inversely $\propto k_{x},k_{y}$.
Since, in the absence of SOC, the overall angular momentum of the spin-1 system is conserved, the corresponding topological charge of the system is $|Q|=0$.
In the presence of SOC, the ground state structures emerge as a result of the interplay
@@
Figure 5: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-sine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
among the contributions from the magnetic field, spin-orbit coupling, and the spinor order parameter. For SOC strength, $\gamma=0.5$, we show the density profiles of the $m_{F}=1,m_{F}=0$ and $m_{F}=-1$ components in Fig. 5 ($a1-a3$). The system experiences an induced anisotropy as a result of rotational symmetry breaking due to SOC. This induces distinctive ground state structures, giving rise to the formation of unique configurations such as vortices, skyrmions, and unconventional spin textures. The vorticity profiles (see Fig. 5 ($b1-b3$)) depicts the existence of vortices, anti-vortices in the spin-1 condensate. Due to the induced anisotropy, the vortices or anti-vortices are not only deformed in shape, but also biased in opposite axes for the $m_{F}=1$ and $m_{F}=-1$ components, respectively. The $m_{F}=0$ component experiences density modulations in both directions as a consequence of the spin-exchange interaction, SOC and external magnetic field.
The k-space density structures (as shown in Fig. 5 ($c1-c3$)) are also anisotropically
@@
Figure 6: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 5. The unit length is the oscillator length $l$.
@@
Figure 7: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the cosine-cosine magnetic field $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
elongated in the preferred direction of high densities, thereby justifying the density modulations, arising due to the combined outcome of SOC and the sine-sine magnetic field. Analogous to the previously discussed situations, here also we can associate a density modulation with wavelength $\lambda_{i}$ for additional momentum peak at $k_{i}\neq 0,\quad i=(x,y)$.
@@
We show for the above discussed case, the topological charge distribution in Fig. 6 ($a$) and associated spin texture in Fig. 6 ($b$). The overall spin texture has both skyrmions and anti-skyrmions along the two orthogonal axes. Fig. 6 ($c-d$) are the local amplification of the spin textures showing a skyrmion and anti-skyrmion, respectively.
3.2.2 Case-\@slowromancapii@ Cosine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\mathrm{cos}x,\mathrm{cos}y)$
Here, we choose the magnetic fields as cosine functions in both directions, i.e., $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with equal magnetic strength $B_{0}=0.25$ in both directions. Analogously, we first attempt to understand the ground state structures due to this magnetic field independent of the SO-coupling. The component-wise density profiles are shown in Fig. 7 ($a1-a3$). Much likely the sine-sine magnetic field case, we observe for this case also, due to the interaction of magnetic fields in both directions, a certain stable saddlepoint structure is formed, owing to which a combination of equal vortices and anti-vortices are being formed in the $m_{F}=1$ and $m_{F}=-1$ components such that the overall conservation of angular momentum is obeyed in the spin-1 system. The spin-exchange interaction along with the effect of magnetic field, yields vortex brights in the $m_{F}=0$ component. These local high densities in the $m_{F}=0$ component compensate for the vortices or anti-vortices in the adjacent two components. The vorticity profiles associated to the component densities are presented in Fig. 7 ($b1-b3$), which clearly explains the angular momentum conservation in the system. The topological excitations arise in the system in the form of exotic vortex-anti-vortex configuration at the saddle points of the external magnetic field. The density modulations due to such exotic structures can be explained by the k-space density profiles (see Fig. 7 ($c1-c3$)). For this case also, apart from the central peak, there exists extra momenta both in $k_{x},k_{y}$ conforming the density modulations with wavlength inversely $\propto$ non-zero momentum peaks in both directions. The topological charge is zero for the non SOC case to respect the angular momentum conservation.
Now, we discuss further the effect of the cosine-cosine magnetic field on the SO-coupled spin-1 BEC. For a weak SOC with strength $\gamma=0.5$, the rotational symmetry of the system breaks giving rise to certain topological excitations in the system. The density profiles of the SO-coupled spin-1 BEC are shown in Fig. 8 ($a1-a3$). Due to the symmetry-breaking, an anisotropy is induced in the system which renders a very unique density distribution to three components of the spin-1 BEC. Trivially, due to the spin-mixing interaction, the $m_{F}=1$ and $m_{F}=-1$ components complement each other, with the presence of vortices in one at the loci of high density in the other component. Moreover, the vortices in one are compensated by the anti-vortices in the other. In the $m_{F}=0$ component, criss-cross vortex and vortex-brights are observed at the loci of local high densities at the $m_{F}=\pm 1$ components, as a consequence of the density modulations due to the combined effect of SOC, external magnetic field and spin-mixing dynamics in the system. However, these vortices in the $m_{F}=0$ component have opposite vorticity in the stable state configuration. The vorticity profiles as shown in Fig. 8 ($b1-b3$) gives a clear picture about the topological excitations in the system. In Fig. 8 ($c1-c3$), we present the k-space densities of the three components of the spin-1 BEC. Here also, apart from the central peak, extra momenta along the $k_{x},k_{y}$ directions are observed. These corresponds to the density modulations. Interestingly, for the $m_{F}=0$ component, the k-space density is modulated obliquely as the local high densities are in that direction.
We present the associated topological charge distribution in Fig. 9 ($a$) and the spin texture is shown in Fig. 9 ($b$). For the diagonal vortices in the $m_{F}=0$ component, as the cores are symmetric, the skyrmion and anti-skyrmion structures are circular. However, for elongated

### v2

@@ III.1 In-plane polarized gradient magnetic field
Figure 1: The ground state densities $n$ and momentum distributions $\tilde{n}$ for the spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC. The first and second columns show the density (a1-a3) and momentum (b1-b3) distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components for the $x$-polarised magnetic field $B=B_{0}x$. The third and fourth columns represent the density (c1-c3) and momentum (d1-d3) distributions for the $y$-polarised magnetic field $B=B_{0}y$. The magnetic field strength $B_{0}=0.25$ for both cases. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length.
In this section, we consider a polarized gradient magnetic field either in $x$ or $y$-direction, i.e., $x$-polarised ${B_{x}=B_{0}x}$ or $y$-polarised ${B_{y}=B_{0}y}$. To ensure that the magnetic field is divergenceless,
@@
Figure 2: The ground state structures for the spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the the $x$-polarised magnetic field $B=B_{0}x$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions $n$, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions $\tilde{n}$ of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
an additional magnetic field along the $z$-direction, i.e., $B_{z}=-B_{0}z$ is required. However, as we ensure a very tight binding along the $z$-direction, the magnetic field along this direction will not affect the quasi-2D condensates. The motivation to explore this case is to underline the impact of such direction-dependent in-plane field on the SO-coupled BEC. To begin with, we first study the case in the absence of SOC. The ground state density distributions $n_{1}=|\psi_{1}|^{2}$, $n_{0}=|\psi_{0}|^{2}$ and $n_{-1}=|\psi_{-1}|^{2}$ for the $x-$polarised magnetic field are shown in Fig. 1 (a1-a3) and that for the $y-$polarised case are shown in Fig. 1 (c1-c3). In the absence of SOC, the density $n_{0}$ is fragmented into two halves about the $x$-axis with both fragments being phase coherent individually, while the densities $n_{1}$ and $n_{-1}$ have a overall disk-shaped condensate with high densities at the $x=0$ line. However, for the $y$-polarised case, the density modulations are observed along the $y=0$ line. Due to the intricate interplay between the spin-$1$ order parameter and the magnetic field bias, the density profiles $n_{\pm 1}$ and $n_{0}$ show reverse features, i.e., the central high density peak is observed in $n_{0}$ while the fragments are observed in $n_{1},n_{-1}$. In the absence of SOC, there are no vortices formed in the system as the polarised gradient magnetic field doesn’t have any saddle-point. However, the momentum space density profiles (as shown in Fig. 1 (b1-b3) and Fig. 1 (d1-d3) for the $x$-polarised and $y$-polarised case, respectively), emphasizes the unique structure of the condensates owing to the direction-dependent density modulations. These density modulations can be associated with the momentum distribution. Along with the central peak at $k_{x},k_{y}=0$, there exists density peaks at some higher $k_{x}$ value for the $x$-polarised magnetic field and similarly, some higher $k_{y}$ value for the $y$-polarised case. These $k_{i}$ values could be associated to a density wave with wavelength $\lambda_{i}=2\pi/k_{i},\quad i=x,y$. In other words, extra peaks in momemtum distribution at $k_{x},k_{y}\neq 0$ corresponds to density modulations in the real space, such that the wavelength $\lambda_{i}$ of this modulation is $\propto 1/k_{i}$.
Now, with SOC in play, our investigation centers to unravel the competition between the SOC and external magnetic field. Depending on the choices of the external magnetic field applied, the ground state structures for the SO-coupled spin-1 BEC are different. We show the SO-coupled ground state structures for the $x$-polarised magnetic field in Fig. 2. The ground state densities $n_{\pm 1},n_{0}$ exhibit modulations along the $x=0$ line as shown in Fig. 2 (a1-a3).
@@
Figure 3: (a) shows the topological charge distribution, (b) shows the spin texture of the system and (c) shows the local amplification of the spin texture, marked by the black square in (b) for the SO-coupled spin-1 BEC shown in Fig. 2. The color of the arrows represent the $S_{z}$ values. Units are dimensionless.
@@
Our numerical simulations show that the formation of vortices in the three components is pertinent to the interplay of SOC and the external magnetic field. As the SOC couples the intrinsic spin to the particles’ motion, it imparts some extra angular momentum to the system which results in the formation of vortices in the system. The SOC influences with the magnetic field in such a way that the effective magnetic field is modified. In addition to this, as there are no saddle-points in the magnetic field, these vortices appear at the $x=0$ line under the influence of SOC. The vorticity profile associated to each component is shown in Fig. 2 (b1-b3). Due to the anisotropic nature of the SOC, the vortices are slightly deformed along the $y$-direction as the component densities are modulated along the $x=0$ line. The k-space density profiles (as shown in Fig. 2 (c1-c3)) depict that apart from the central peaks, some extra density peaks predominantly along the $k_{x}$-direction are observed. These extra peaks at $k_{x}\neq 0$ corresponds to the density modulations along the $x$-direction. This is due to the fact that SOC influences the distribution of particles in the k-space, similar to the impact of SOC as seen in the distribution of electronic states as a function of momentum in solid-state physics [107].
@@
The combined effect of the in-plane direction dependent gradient field and the SOC yield intricate spin textures in the spin-1 BEC. The topological charge distribution $q(\mathbf{r})$ and the corresponding spin texture is shown in Fig. 3 (a) and Fig. 3 (b), respectively. We can identify the presence of a stable spin configuration characterized by a converging pattern centred at a point carrying +ve charge, called skyrmion, from the spin texture. The local amplification of the skyrmion spin texture is shown in Fig. 3 (c). The spin texture is elongated along the $x=0$ line due to the SOC-induced anisotropy in the system.
@@
Figure 4: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the sine-sine magnetic field $B_{x}=B_{0}\sin x$ and $B_{y}=B_{0}\sin y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
For the SO-coupled spin-1 BEC in $y$-polarised gradient magnetic field, we can analogously understand the density modulations along the $y=0$ line for all the three components. However, these density modulations converge into the formation of anti-vortices in the system, and the corresponding spin textures have anti-skyrmions along the $x$-axis. It is interesting to note that the state which had fragments in the absence of SOC, now has more vortices or anti-vortices as we observe for the case with SOC. The number of vortices increases as we increase the SOC strength $\gamma$ (see appendix A). However, for the polarised magnetic field, they appear in the same alignment, with the vortex spacing being reduced and more deformed.
@@
Figure 5: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-sine magnetic field $B_{x}=B_{0}\sin x$ and $B_{y}=B_{0}\sin y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
III.2 Cross-polarised sinusoidally varying in-plane magnetic fields
Thus far, we have discussed how impactful the polarised gradient field can be on the ground state structure of a SO-coupled spin-1 BEC. To further exploit this direction-dependent interplay of external magnetic field and SOC, we now choose sinusoidally varying in-plane magnetic fields in both $x$- and $y$-directions. In general, to make the magnetic field divergenceless, one has to apply a magnetic field $B_{z}=-B_{0}zf(x,y)$ in the $z$-direction, where $f(x,y)$ are sinusoidal functions in the $x$-$y$ plane. However, as we consider a quasi-2D geometry, for tight binding along the $z-$direction, $B_{z}$ will not affect the spinor condensate. In this section, we center our focus only on the in-plane sinusoidal magnetic field. Experimentally, a polarised sinusoidal field can be realised using a similar technique discussed in [24]. Campbell et al., $2016$ has also demonstrated a similar in-plane sinusoidal field in their work [93]. We discuss below various combinations of $B(\mathbf{r})=(B_{x},B_{y})$ in different cases.
@@ III.2 Cross-polarised sinusoidally varying in-plane magnetic fields
Figure 6: (a) shows the topological charge distribution, (b) shows the spin texture of the system and (c),(d) shows the local amplification of the spin texture, marked by green and black squares, respectively, for the SO-coupled spin-1 BEC shown in Fig. 5. The color of the arrows represent the $S_{z}$ values. Units are dimensionless.
III.2.1 Case-\@slowromancapi@: Sine-Sine magnetic field ${B}(\mathbf{r})=B_{0}(\sin x,\sin y)$
For the present case, we choose the magnetic fields as sine functions in both directions, i.e., ${B_{x}}=B_{0}\sin x$ and ${B_{y}}=B_{0}\sin y$. We first present the ground state structures for the no SOC case in Fig. 4. The density profiles $n_{1},n_{0}$ and $n_{-1}$ of the condensates $m_{F}=1,m_{F}=0$ and $m_{F}=-1$, respectively are shown in Fig. 4 (a1-a3). The sine-sine type of magnetic field results in the density modulations in $m_{F}=\pm 1$ components with the appearance of topological exciations in the form of a triad of alternating vortex-anti-vortex along both $x$- and $y$-axes. This is due to the fact that the applied magnetic fields from both directions generate certain saddle-point structures such that the nature of the triads in $m_{F}=1$ and $m_{F}=-1$ components are complementary to each other. The contrasting nature of the triads is due to the fact that the magnetic field appears as conjugate for the $m_{F}=1$ and $m_{F}=-1$ components in the GP equations (II-II). Furthermore, combining with the spin-mixing term, the magnetic field yields vortex brights in the $m_{F}=0$ component at the loci (saddle-points of $B(\rm{r})$) of vortices in the $m_{F}=\pm 1$ components so as to conserve the overall angular momentum of the spin-1 BEC in the absence of SOC. Thus, the $m_{F}=0$ component has no vortices, instead a high density peak is observed at the center of the condensate. The vorticity profiles associated to the condensate densities are depicted in Fig. 4 (b1-b3). The k-space density profiles (see Fig. 4 (c1-c3)) support observed density modulations in the system. For the $m_{F}=1$ and $m_{F}=-1$ components, apart from the central peak, we observe some additional density peaks both in the $k_{x}$ and $k_{y}$ directions which correspond to the density modulations with wavelength inversely $\propto k_{x},k_{y}$. In the absence of SOC, the total angular momentum of the spin-1 system remains conserved, resulting in a topological charge of $|Q|=0$. Consequently, vortices or anti-vortices within the system do not play an active role in shaping the spin texture.
@@ III.2.1 Case-\@slowromancapi@: Sine-Sine magnetic field ${B}(\mathbf{r})=B_{0}(\sin x,\sin y)$
Figure 7: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the cosine-cosine magnetic field $B_{x}=B_{0}\cos x$ and $B_{y}=B_{0}\cos y$ with strength $B_{0}=0.25$. The first column (a1-a3) shows the density distributions, second column (b1-b3) shows the vorticity profiles and the third column (c1-c3) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The densities $n$ and $\tilde{n}$ are expressed in units of $l^{-2}$ and $l^{2}$, respectively, where $l=2.41\mu m$ is the characteristic length. The vorticity $\Omega$ is expressed in dimensionless units.
In the presence of SOC, the ground state structures emerge as a result of the interplay among the contributions from the magnetic field, spin-orbit coupling, and the spin-mixing dynamics. We show the density profiles of the $m_{F}=1,m_{F}=0$ and $m_{F}=-1$ components in Fig. 5 (a1-a3). The system experiences an induced anisotropy as a result of rotational symmetry breaking due to SOC. The vorticity profiles (see Fig. 5 (b1-b3)) depict the existence of vortices, anti-vortices in the spin-1 condensate. Due to the induced anisotropy, the vortices or anti-vortices are not only deformed in shape, but also are biased in opposite axes $x$ and $y$, for the $m_{F}=1$ and $m_{F}=-1$ components, respectively. In other words, for this case the vortices and anti-vortices are respectively nucleated at the two orthogonal axes. The $m_{F}=0$ component experiences density modulations in both directions as a consequence of the spin-exchange interaction, SOC and external magnetic field. The k-space density structures (as shown in Fig.5 (c1-c3)) have extra peaks in the preferred direction of high spatial densities, in addition to the anisotropically elongated central peak in the $m_{F}=\pm 1$ components. Analogous to the previously discussed situations, here also we can associate a density modulation, arising due to the combined outcome of SOC and the sine-sine magnetic field, with wavelength $\lambda_{i}$ for additional momentum peak at $k_{i}\neq 0,\quad i=(x,y)$.
@@
We show for the above discussed case, the topological charge distribution in Fig. 6 (a) and associated spin texture in Fig. 6 (b). The overall spin texture has both anti-skyrmions  [^1] Here, we characterise anti-skyrmions on the basis of their $S_{z}$ values and the -ve topological charge they carry. and skyrmions along the two orthogonal axes $x$ and $y$, respectively. Fig. 6 (c-d) are the local amplification of the spin textures showing a skyrmion and anti-skyrmion, respectively. The spin textures arise in the system due to the SOC, however the arrangement of various spin structures depend on the choice of external in-plane magnetic field.